OpenStudy (anonymous):

find all solutions in the internal (0.2pi] cos 2x -square 2 sin-1 help

OpenStudy (anonymous):

it appears that your answer will be a number set. so let x=0.2 and solve, then let x = pi and solve. with the two solutions you get, these will be the number set representing the bounds of 'all solutions' message me if you get stuck ^_^

OpenStudy (anonymous):

i'm stuck :/

OpenStudy (anonymous):

Help :/

OpenStudy (anonymous):

(0.2pi] means (0.2, pi] right?

OpenStudy (anonymous):

this cos 2x -square 2 sin-1 means this cos (2x) - sin^2 (-1) right?

OpenStudy (anonymous):

^_^ LOL

OpenStudy (anonymous):

\[\cos2x - \sqrt{2} \sin-1\]

zepdrix (zepdrix):

\[\Large \cos2x - \sqrt{2} \sin-1=0\]Recall your `Cosine Double Angle Identity`:\[\Large \color{#CC0033}{\cos2x\quad=\quad 1-2\sin^2x}\]

zepdrix (zepdrix):

\[\Large 1-2\sin^2x-\sqrt2 \sin x-1=0\]Hmm looks like the 1's cancel out, that's convenient. Do you need help from this point?

OpenStudy (anonymous):

yes i'm

zepdrix (zepdrix):

\[\Large \cancel1-2\sin^2x-\sqrt2 \sin x\cancel{-1}=0\]Multiply each side by -1 giving us,\[\Large 2\sin^2x+\sqrt2 \sin x\quad=\quad 0\]

zepdrix (zepdrix):

Factor a sinx out of each term,\[\Large \sin x(2\sin x+\sqrt2)\quad=\quad 0\]To continue, set each factor equal to zero, then solve for sinx in each case,\[\Large \sin x=0\qquad\qquad\qquad 2\sin x+\sqrt2=0\]

zepdrix (zepdrix):

Hmm you should get some nice special angles in each case.

OpenStudy (anonymous):

like you

OpenStudy (anonymous):

0, \[-\sqrt{2}/2\]

OpenStudy (anonymous):

right